OpenStudy (anonymous):

evaluate: 1.) (-3y^5) (-7y^4) 2.) 5a^-8 b^4 3.) pq^-3 r^-4 / 3u^-5 4.) 14x^-2 yz^-4 / 2xyz

OpenStudy (igreen):

You have to solve all of these? @lyssaaaa

OpenStudy (anonymous):

yep it's actually up to ten but nah I think I can do the others

OpenStudy (igreen):

(-3y^5)(-7y^4) For the first one, just multiply -3 and -7, then add the exponents.

OpenStudy (igreen):

That will give you 21y^9

OpenStudy (anonymous):

ohhhh

OpenStudy (igreen):

Is the 2nd one multiplication?

OpenStudy (anonymous):

I think so. It says 5a raised to -8 b raised to 4 that's it

OpenStudy (igreen):

Oh, do you have choices? Because you can't simplify that anymore..the same goes with the third one..you can't simplify it anymore.

OpenStudy (igreen):

In the fourth one: 14x^-2 yz^-4 / 2xyz Are 14x^-2 and yz^-4 being multiplied?

OpenStudy (anonymous):

It doesn't have choices

OpenStudy (igreen):

Well, just say #2 and #3 are already simplified.

OpenStudy (anonymous):

I think I have to simplify it first ?

OpenStudy (anonymous):

yep okay how abt num 4

OpenStudy (anonymous):

isn't it like 14x^-2 simplify first with 2x?

OpenStudy (igreen):

For the fourth one, just divide 14 by 2, and subtract 1 for every exponent in the numerator. 14x^-2 yz^-4 / 2xyz That will give us: 7x^-3 z^-5 It cancels out the y-variable.

OpenStudy (anonymous):

Ohhh

OpenStudy (igreen):

Yes, you would simplify the 14x^-2 first.

OpenStudy (anonymous):

I still have a few questions is it okay ?

OpenStudy (igreen):

For the third one, you can't divide because the denominator doesn't have a similar variable with the numerator.

OpenStudy (igreen):

Yeah, sure.

OpenStudy (anonymous):

What abt (5m^-1 n^2) (-3m^3 n)

OpenStudy (igreen):

\((5m^{-1} n^2) (-3m^3 n)\) Multiply the bases, add the exponents. \(5 * -3 = -15.\) \(m^{-1} + m^3 = m^2\) \(n^2 + n = n^3\) So: \(-15m^2n^3\)

OpenStudy (anonymous):

Ohhh hb (-3x^-2 y)^-1 u will distribute -1 right?

OpenStudy (anonymous):

And last 10^-5 / 10^6 = 10 am I right?

OpenStudy (anonymous):

last question promise (-5)^-2

OpenStudy (igreen):

(-5)^-2 = -0.04